Find in each case the number or numbers that can replace k so that the resulting ordered number pair will be on the graph in the given equation.
x + 2y = 5;(k,2)

x + 3y = 10;(13,k)

x - y = 0;(k,k)

x + y = k;(2,5)

Thank you so much for you help!

Answer by rfer(16322)

A)
x+3*2=5
x+6=5
x=5-6
x=-1
(-1,2)
B)
13+3y=10
3y=10-13
3y=-3
y=-3/3
y=-1
(13,-11)